OpenStudy (anonymous):

Find a pattern for each sequence. Use the pattern to identify the next two terms or figures. 1. 3, -6, 18, -72, 360 -500; 10,600 -2,160; 15,120 -4,320; 20,113 2,160; 15,120

OpenStudy (amistre64):

might be geometeric since it alternates signs like that

OpenStudy (amistre64):

or it could be some random variation .... hmmm

OpenStudy (amistre64):

3(-2) = -6 -6(-3) = 18 18(-4) = -72 ... that seems to work

OpenStudy (anonymous):

so whats the pattern?

OpenStudy (amistre64):

A{n} = A{n-1}*(-1)^n*(n+1) looks plausible for an equation to solve? the easiest thing to do is just brute math it once you see the pattern

OpenStudy (anonymous):

still typeing? lol :)

OpenStudy (amistre64):

A{n} = -(n) *A{n-1} A{2} = -(2) * (3) = -6 A{3} = -(3) * (-6) = 18 A{4} = -(4) * (18) = -72 trying to see whats in the question ☺

OpenStudy (anonymous):

ok ill wait :)

OpenStudy (amistre64):

3, -6, 18, -72, 360 -500; <- whats the semicolon for? 10,600 -2,160; <- are these different sequences? 15,120 -4,320; 20,113 2,160; 15,120

OpenStudy (amistre64):

i think I see it now; those are the answer choices right ☺

OpenStudy (anonymous):

yes :)

OpenStudy (amistre64):

Q1 3, -6, 18, -72, 360 a) -500; 10,600 b) -2,160; 15,120 c) -4,320; 20,113 d) 2,160; 15,120 ...formating helps to understand the question ;)

OpenStudy (anonymous):

so how would i do that?

OpenStudy (amistre64):

we know the next one is - so d is out

OpenStudy (amistre64):

format? just hit enter while typing the question the enter key doesnt "post"; you actually have to click or tab over to the post

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

we want 6 and 7 A{n} = -(n) * A{n-1} A{6} = -(6) * 360 = -2160 and the only option for that is 'b'; but lets carry it out to the end A{7} = -(7) * -2160 = 15120

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

I was going to try to find a particular equation to solve; but since it was so small, i figured...nah lol

OpenStudy (anonymous):

lol :D thanks soo much!